OpenStudy (anonymous):

use substitution to solve: 3x - 2y = 9 and 2x + 5y = 19

OpenStudy (poofypenguin):

Solve one of the equations for x and then substitute this equation into the other one.

OpenStudy (poofypenguin):

So solving 3x - 2y = 9 for x we get: x=(9+2y)/3

OpenStudy (poofypenguin):

then we substitute x=(9+2y)/3 into 2x + 5y = 19 in order to solve for y: 2[(9+2y)/3] + 5y = 19 [(18+4y)/3]+5y=19 18+4y=(19-5y)*3 18+4y=57-15y 15y+4y=57-18 19y=39 y=39/19

OpenStudy (poofypenguin):

Now we can substitute y=39/19 into x=(9+2y)/3 to get the value of x: x=(9+2y)/3 x=[9+2(39/19)]/3 x=83/19